# Conventional and practical metallic superconductivity arising from repulsive Coulomb coupling

### v1

@@ V Summary and Conclusion
We conclude by asserting that the case for conventional superconductivity in metals with an experimentally observable $T_{c}$ arising from plasmon exchange (or equivalent mechanisms associated with electron-electron interactions) has not been established at all in spite of there being many claims. Our work shows that Coulomb coupling induced (or plasmon-induced) conventional $s$-wave superconductivity in normal metals to be unlikely.
@@ Acknowledgment
The authors thankfully acknowledge helpful communications with Steve Kivelson, Shiwei Zhang and Shuyang Wang. This work is supported by the Laboratory for Physical Sciences. J.S. acknowledges support from the Joint Quantum Institute.

### v2

@@ V Summary and Conclusion
We conclude by asserting that the case for conventional superconductivity in metals with an experimentally observable $T_{c}$ arising from plasmon exchange (or equivalent mechanisms associated with electron-electron interactions) has not been established at all in spite of there being many claims. Our work shows that Coulomb coupling induced (or plasmon-induced) conventional $s$-wave superconductivity in normal metals to be unlikely. Although we cannot decisively rule out the possibility of weak low-$T_{c}$ SC arising from Coulomb interactions (since the problem is not perturbatively amenable because of the inapplicability of the Migdal theorem), we can certainly rule out claims of plasmon-mediated SC with effectively high $T_{c}$, and our work establishes the strict guidelines that future theories of this problem must obey.
@@ Acknowledgment
The authors thankfully acknowledge helpful communications with Steve Kivelson and Shiwei Zhang. This work is supported by the Laboratory for Physical Sciences. J.S. acknowledges support from the Joint Quantum Institute.
Appendix A Details on the screened Coulomb interaction in 2D
Similar to the 3D case considered in Sec. II, here we consider a two-dimensional electron gas with dispersion $\varepsilon_{\mathbf{p}}=p^{2}/2m$ and spin degeneracy 2. At zero temperature, the Lindhard polarizability
$\chi(\mathbf{q},iq_{n})=2\int\frac{d\mathbf{p}}{(2\pi)^{2}}\frac{f_{\mathbf{p}}-f_{\mathbf{p}+\mathbf{q}}}{iq_{n}+\varepsilon_{\mathbf{p}}-\varepsilon_{\mathbf{p}+\mathbf{q}}},$ (48)
can be calculated analytically ($f_{\mathbf{p}}=\theta(\varepsilon_{\mathbf{p}}-\mu)$). Substitute $iq_{n}=\omega+i\epsilon,\epsilon\to 0$ for the retarded response and define dimensionless variables $\tilde{q}=q/2k_{F}$, $\tilde{\omega}=\omega/4\varepsilon_{F}$ and $\tilde{\chi}=\chi/d(\varepsilon_{F})$ (where $d(\varepsilon_{F})$ is the density of states at the Fermi energy $\varepsilon_{F}$, and $k_{F}$ is the corresponding Fermi momentum), we have [Stern1967, Ando1982]
$\tilde{\chi}(\tilde{q},\tilde{\omega}+i\epsilon)=\frac{1}{2\tilde{q}}\Bigg[-i\sqrt{1-\left(\tilde{q}-\frac{\tilde{\omega}+i\epsilon}{\tilde{q}}\right)^{2}}\\ +i\sqrt{1-\left(\tilde{q}+\frac{\tilde{\omega}+i\epsilon}{\tilde{q}}\right)^{2}}\Bigg]-1,$ (49)
where $\sqrt{1-(x\pm i\epsilon)^{2}}$ for real $x$ goes to $\sqrt{1-x^{2}}$ for $|x|<1$ and $\mp i\operatorname{sgn}(x)\sqrt{x^{2}-1}$ for $|x|>1$.
Under RPA, the dielectric function is approximated by
where $v_{c}(q)=2\pi e^{2}/q$ is the bare 2D Coulomb coupling. The screened Coulomb interaction is
Defining the dimensionless interaction $u=\frac{\pi e^{2}}{k_{F}}\tilde{u}$, we can express everything using dimensionless parameters as
$\tilde{u}_{\text{RPA}}(\tilde{q},\tilde{\omega})=\frac{1}{\tilde{q}-\frac{r_{s}}{\sqrt{2}}\tilde{\chi}(\tilde{q},\tilde{\omega}+i\epsilon)},$ (52)
where $r_{s}=(\pi a_{0}^{2}n)^{-1/2}$ and $n$ is the 2D electron density. The function is plotted in Fig. 5.
For PPA we have [Stern1967, Ando1982]
$\varepsilon_{\text{PPA}}(\mathbf{q},\omega)$ $=1-\frac{1}{\omega^{2}}\left(\frac{2\pi ne^{2}q}{m}+\frac{3}{4}q^{2}v_{F}^{2}\right)$ (53)
$=1-\frac{1}{\tilde{\omega}^{2}}\left(\frac{r_{s}}{\sqrt{8}}\tilde{q}+\frac{3}{4}\tilde{q}^{2}\right),$
The resulting Coulomb interaction is plotted in Fig. 6.